# The Rees algebra and analytic spread of a divisorial filtration

### v1

@@ Abstract.
In this paper we investigate some properties of Rees algebras of divisorial filtrations and their analytic spread. A classical theorem of McAdam shows that the analytic spread of an ideal $I$ in a formally equidimensional local ring is equal to the dimension of the ring if and only if the maximal ideal is an associated prime of $R/\overline{I^{n}}$ for some $n$. We show in Theorem 1.6 that McAdam’s theorem holds for ${\mathbb{Q}}$-divisorial filtrations in an equidimensional local ring which is essentially of finite type over a field. This generalizes an earlier result for ${\mathbb{Q}}$-divisorial filtrations in an equicharacteristic zero excellent local domain by the author. This theorem does not hold for more general filtrations.
@@ 1.1. Graded Filtrations, Rees Algebras and Analytic Spread
We always have ([13, Lemma 3.6]) that
@@
so the classical inequality $\ell(I)\leq\dim R$ of ideals (Proposition 5.1.6 [36]) extends to arbitrary graded filtrations.
@@
Let $k$ be a field and $R$ be the localization of a polynomial ring $R=k[x,y,z]_{(x,y,z)}$. Given any function $\sigma:{\mathbb{Z}}_{>0}\rightarrow{\mathbb{Z}}_{>0}$, define the graded filtration =$\mathcal{I}=\{I_{n}\}$ by $I_{0}=R$ and $I_{n}=(z^{n+2})+z^{n+1}(x,y)^{\sigma(n)}$ for $n\geq 1$. Then
@@
1.2. McAdam’s theorem for divisorial fltrations
@@
([29], [36, Theorem 5.4.6]) Let $R$ be a formally equidimensional local ring and $I$ be an ideal in $R$. Then $m_{R}\in\mbox{Ass}(R/\overline{I^{n}})$ for some $n$ if and only if $\ell(I)=\dim(R)$.
The assumption of being formally equidimensional is not required for the if direction of Theorem 1.2 (this is Burch’s theorem, [4], Proposition 5.4.7 [36]).
@@
This follows from [34] or the proof of Theorem 10.2.2 [36], Proposition 10.4.3 [36] and Scholie IV.7.8.3 (x) [20]. ∎
@@
The following theorem is proven in Section 3.2. Theorem 1.4 is proven for excellent equicharacteristic zero local domains in Theorem 1.3 [11].
@@
Let $R$ be an equidimensional local ring which is essentially of finite type over a field and let =$\mathcal{I}=\{I_{n}\}$ be a ${\mathbb{Q}}$-divisorial filtration on $R$. If $m_{R}\in\mbox{Ass}(R/I_{m_{0}})$ for some $m_{0}\in{\mathbb{Z}}_{>0}$, then the analytic spread of $\mathcal{I}$ is $\ell(\mathcal{I})=\dim R$.
The conclusions of Theorem 1.4 do not hold for more general filtrations. In fact, there exist ${\mathbb{R}}$-divisorial filtrations $\mathcal{I}$ on an excellent equicharacteristic zero local domain such that the conclusions of Theorem 1.4 are false, as shown in Example 1.5 [11].
The following theorem is proven for domains in [13]. The proof is the same for local rings, using the notation on valuations of a ring and divisorial filtrations given in Section 2.2.
Theorem 1.5.
(Theorem 1.4 [13] extended to local rings) Suppose that $R$ is a local ring and =$\mathcal{I}=\{I_{n}\}$ is an ${\mathbb{R}}$-divisorial filtration on $R$. Suppose that $\ell(\mathcal{I})=\dim R$. Then there exists a positive integer $n_{0}$ such that $m_{R}$ is an associated prime of $I_{n}=\overline{I_{n}}$ for all $n\geq n_{0}$.
@@
Let $R$ be an equidimensional local ring which is essentially of finite type over a field. Let =$\mathcal{I}=\{I_{n}\}$ be a ${\mathbb{Q}}$-divisorial filtration on $R$. Then the following are equivalent.
@@
3) implies 1) of Theorem 1.6 is Theorem 1.4 and 1) implies 2) follows from Theorem 1.5.
Theorem 1.6 is proven for excellent equicharacteristic zero local domains in Theorem 1.4 [11].
@@ 1.3. Hilbert functions of divisorial $m_{R}$-filtrations
In this subsection we consider the Hilbert function of a ${\mathbb{Q}}$-divisorial filtration =$\mathcal{I}=\{I_{n}\}$ of $m_{R}$-primary ideals on a $d$-dimensional normal excellent local ring $R$. We investigate asymptotic properties of the function
@@
A good property of this function is that the multiplicity
@@
always exists. In fact, this limit always exists for arbitrary graded families of $m_{R}$-primary ideals in analytically unramified local rings ([6, Theorem 1.1] and [7, Theorem 4.2]). This theorem builds on earlier work by [18], [30], [31], [23] and [26].
We consider the asymptotic properties of the difference function $n\mapsto\ell_{R}(I_{n}/I_{n+1})$. In the case of an $I$-adic filtration, this function is a polynomial of degree $d-1$ for $n\gg 0$. We always have that
(4) $\limsup_{n\rightarrow\infty}\frac{\lambda_{R}(I(n)/I(n+1))}{n^{d-1}}$
is finite for ${\mathbb{Q}}$-divisorial filtrations of $m_{R}$-primary ideals, as is shown in Lemma 4.1. However, there are examples of graded families of ideals such that this limsup is infinite (Theorem 4.3 [7]). Even when this limsup is finite, it may be that the limsup is not a limit for graded filtrations of $m_{R}$-primary ideals (Theorem 4.6 [7]).
The function (3) is very nice in dimension 2. In Section 8 [14], it is shown that if $R$ is a two dimensional normal excellent local ring. and =$\mathcal{I}=\{I_{n}\}$ is a ${\mathbb{Q}}$-divisorial filtration of $m_{R}$-primary ideals on $R$, then there exists a quadratic polynomial $P(n)$ with rational coefficients, a function $f(n)$ and a constant $c$ such that
@@
with $|f(n)|<c$ for all $n$. In Theorem 9 [14], it is shown that if $R$ has equicharacteristic zero then $f(n)$ is eventually periodic. Example 5 [14] shows that $f(n)$ may not be eventually periodic if $R$ has characteristic $p>0$.
@@
(5) $n\mapsto\lim_{n\rightarrow\infty}\frac{\lambda_{R}(I_{n}/I_{n+1})}{n^{d-1}}$
@@
As we have seen above, this limit does always exist and is even a rational number when the dimension of $R$ is $d=2$.
@@
It is shown in the formula before Lemma 2.2 [9] and Theorem 4.1 in [9] that
@@
where $D_{n}=\lceil n\alpha\rceil\overline{S}+nF$, $\alpha=\frac{3}{9-\sqrt{3}}$. It is further shown that the divisors $-D_{n}$ are nef divisors on $U$. Thus by formula (9) and Proposition 2.4 of [9],
@@ 1.4. Symbolic powers with infinitely many Rees valuations
If $I$ is an ideal in an excellent local ring, then we have seen in Theorem 1.3 that the filtration of powers of integral closures $\{\overline{I^{n}}\}$ of the powers of $I$ is a ${\mathbb{Z}}$-divisorial filtration. In particular, the set of all Rees valuations of all of the ideals $I^{m}$ over all $m$ is a finite set. It is just the set of Rees valuations of $I$ itself. In Section 5 we consider the question of if this set is finite for the ideals in a ${\mathbb{Z}}$-divisorial filtration; that is, is the set of all Rees valuations of the ideals $I_{n}$ finite when =$\mathcal{I}=\{I_{n}\}$ is a ${\mathbb{Z}}$-divisorial filtration? We show this is not the case. In Section 5 we construct an example of a height one prime ideal in a two dimensional normal local ring $R$ such that for all $n\geq 1$, the symbolic power $P^{(n)}$ has a Rees valuation other then the $PR_{P}$-adic valuation $v_{P}$, but for $1\leq m<n$, $v_{P}$ is the only common Rees valuation of $P^{(m)}$ and $P^{(n)}$. In particular, the set of Rees valuations of $P^{(n)}$ over all $n\geq 1$ is infinite.
@@ 1.5. Acknowledgements
The author would like to thank Hailong Dao and Jonathan Montaño for discussions and questions about Hilbert polynomials of filtrations and to thank Linquin Ma, Thomas Polstra and Karl Schwede for a discussion about the cardinality (finite or infinite) of the set of Rees valuations of the symbolic powers of a prime ideal. He thanks Suprajo Das for suggesting the use of alterations in the generalization to arbitrary ground fields of the assumption that the ground field be of characteristic zero in Theorem 1.3 of [11]

### v2

@@ Abstract.
In this paper we investigate some properties of Rees algebras of divisorial filtrations and their analytic spread. A classical theorem of McAdam shows that the analytic spread of an ideal $I$ in a formally equidimensional local ring is equal to the dimension of the ring if and only if the maximal ideal is an associated prime of $R/\overline{I^{n}}$ for some $n$. We show in Theorem 1.5 that McAdam’s theorem holds for ${\mathbb{Q}}$-divisorial filtrations in an equidimensional local ring which is essentially of finite type over an excellent local ring of dimension less than or equal to 3. This generalizes an earlier result for ${\mathbb{Q}}$-divisorial filtrations in an equicharacteristic zero excellent local domain by the author. This theorem does not hold for more general filtrations.
@@ 1.1. Graded Filtrations, Rees Algebras and Analytic Spread
We always have ([14, Lemma 3.6]) that
@@
so the classical inequality $\ell(I)\leq\dim R$ of ideals( [39, Proposition 5.1.6]) extends to arbitrary graded filtrations.
@@
Let $R=k[x,y,z]_{(x,y,z)}$ be the localization of a polynomial ring over a field $k$. Given any function $\sigma:{\mathbb{Z}}_{>0}\rightarrow{\mathbb{Z}}_{>0}$, define the graded filtration =$\mathcal{I}=\{I_{n}\}$ by $I_{0}=R$ and $I_{n}=(z^{n+2})+z^{n+1}(x,y)^{\sigma(n)}$ for $n\geq 1$. Then
@@
1.2. McAdam’s theorem for divisorial filtrations
@@
([31], [39, Theorem 5.4.6]) Let $R$ be a formally equidimensional local ring and $I$ be an ideal in $R$. Then $m_{R}\in\mbox{Ass}(R/\overline{I^{n}})$ for some $n$ if and only if $\ell(I)=\dim(R)$.
The assumption of being formally equidimensional is not required for the if direction of Theorem 1.2 (this is Burch’s theorem, [4], [39, Proposition 5.4.7 ]).
@@
This follows from [36] or the proof of [39, Theorem 10.2.2], [39, Proposition 10.4.3] and [21, Scholie IV.7.8.3 (x)]. ∎
@@
The following theorem is proven for local domains in [14]. The proof is the same for local rings, using the notation on valuations of a ring and divisorial filtrations given in Section 2.2.
@@
(Theorem 1.4 [14] extended to local rings) Suppose that $R$ is a local ring and =$\mathcal{I}=\{I_{n}\}$ is an ${\mathbb{R}}$-divisorial filtration on $R$. Then
  1. a)
    $I_{n}=\overline{I_{n}}$ for all $n>0$ and
  2. b)
    Suppose that $\ell(\mathcal{I})=\dim R$. Then there exists a positive integer $n_{0}$ such that $m_{R}$ is an associated prime of $I_{n}$ for all $n\geq n_{0}$.
@@
Let $R$ be an equidimensional local ring which is essentially of finite type over an excellent local ring of dimension less than or equal to 3. Let =$\mathcal{I}=\{I_{n}\}$ be a ${\mathbb{Q}}$-divisorial filtration on $R$. Then the following are equivalent.
@@
3) implies 1) of Theorem 1.5 is Theorem 3.5 and Temkin’s theorem on existence of regular alterations [41, Theorem 1.2.5]. Temkin’s theorem shows that if $X$ is an integral scheme which is of finite type over an excellent domain of dimension less than or equal to three, then there exists a projective regular alteration $X^{\prime}\rightarrow X$. Alterations and regular alterations are defined in Definition 3.4 and just after this definition. 1) implies 2) follows from Theorem 1.4.
Theorem 1.5 is proven for excellent equicharacteristic zero local domains in [12, Theorem 1.4], using the existence of resolutions of singularities of excellent equicharacteristic zero domains ([23] or [40]). The proof in [12] shows that Theorem 1.5 is true for excellent local domains of dimension $\leq 3$, using the existence of resolutions of singularities of excellent domains of dimension $\leq 3$ proven in [6].
The conclusions of 3) implies 1) of Theorem 1.5 do not hold for more general filtrations. In fact, there exist ${\mathbb{R}}$-divisorial filtrations $\mathcal{I}$ on an excellent equicharacteristic zero local domain such that the conclusions of Theorem 3.5 are false, as shown in Example 1.5 [12].
The existence of regular alterations over a local domain which is essentially of finite type over a field was first proven by de Jong in [24, Theorem 4.1].
Remark 1.6.
The conclusions of Theorem 1.5 hold whenever $R$ is an equidimensional excellent local ring which has the property that if $P$ is a minimal prime of $R$ and $X\rightarrow\mbox{Spec}(R/P)$ is a birational projective morphism with $X$ integral, then there exists a regular alteration $X^{\prime}\rightarrow X$; in fact, the above proof of Theorem 1.5 (replacing the reference to Temkin’s theorem with this assumption) proves this result.
@@ 1.3. Hilbert functions of divisorial $m_{R}$-filtrations
If $I$ is an ideal in a Noetherian local ring $R$, the function $\lambda_{R}(I^{n}/m_{R}I^{n})$ is a polynomial for large $n$, whose degree is the analytic spread $\ell(I)$ of $I$. The computation of asymptotic properties of the function $\lambda_{R}(I_{n}/m_{R}I_{n})$ is difficult for a divisorial filtration =$\mathcal{I}=\{I_{n}\}$. In the case that $R$ is normal and excellent of dimension two, we show in [11, Theorem 1.6] that for a divisorial filtration $\mathcal{I}$ of $R$, there exists $\alpha\in{\mathbb{Q}}_{\geq 0}$ such that
$\lambda_{R}(I_{n}/m_{R}I_{n})=n\alpha+\sigma(n)$
for $n>0$, where $\sigma(n)$ is a bounded function of $n$. We have that $\alpha>0$ if and only if $\ell(\mathcal{I})=2=\dim R$. The proof uses Zariski decomposition, which does not exist in general if $\dim R\geq 3$.
Now suppose that $\mathcal{I}$ is a divisorial filtration of $m_{R}$ ideals. Then, related to the function $\lambda_{R}(I_{n}/m_{R}I_{n})$, we have the functions
@@
and its first difference $n\mapsto\lambda_{R}(I_{n}/I_{n+1})$. A good property of the function (3) is that the multiplicity
@@
always exists. In fact, this limit always exists for arbitrary graded families of $m_{R}$-primary ideals in analytically unramified local rings ([7, Theorem 1.1] and [8, Theorem 4.2]). This theorem builds on earlier work by [19], [32], [33], [25] and [28].
The multiplicity $e(\mathcal{I})$ is related to the analytic spread $\ell(\mathcal{I})$ as follows.
Lemma 1.7.
Suppose that $R$ is a $d$-dimensional excellent normal local ring and =$\mathcal{I}=\{I_{n}\}$ is a divisorial filtration of $m_{R}$-primary ideals of $R$. Then $e(\mathcal{I})>0$ and $\ell(\mathcal{I})=d$.
We give an outline of the proof. $\ell(\mathcal{I})=d$ by Theorem 1.5. There exist divisorial valuations $w_{j}$ on $R$ (the Rees valuations of $m_{R}$) such that $\mbox{center}_{R}(w_{j})=m_{R}$ for all $j$, and positive integers $b_{j}$ such that
$\overline{m_{R}^{n}}=\cap_{j}I(w_{j})_{mb_{j}}\mbox{ for all $n\geq 0$}.$
Writing
$I_{n}=\cap_{j}I(v_{i})_{na_{i}}$
for suitable divisorial valuations $v_{i}$ with $\mbox{center}_{R}(v_{i})=m_{R}$ and $a_{i}\in{\mathbb{Q}}$, we see by Izumi’s Theorem ([37]) that there exists a positive integer $\beta$ such that $I_{\beta n}\subset\overline{m_{R}^{n}}$ for all $n>0$. Now the ring $\sum_{n\geq 0}\overline{m_{R}^{n}}t^{n}$ is a finite $R[m_{R}t]$-module by [39, Proposition 5.2.1] and [39, Corollary 9.2.1], so
$0<\lim_{n\rightarrow\infty}\frac{\lambda_{R}(R/\overline{m_{R}^{n}})}{n^{d}}\leq\lim_{n\rightarrow\infty}\frac{\lambda_{R}(R/I_{n})}{n^{d}}=e(\mathcal{I}).$
In this subsection, we consider the asymptotic properties of the difference function $n\mapsto\ell_{R}(I_{n}/I_{n+1})$. In the case of an $I$-adic filtration, this function is a polynomial of degree $d-1$ for $n\gg 0$. We always have that
(4) $\limsup_{n\rightarrow\infty}\frac{\lambda_{R}(I_{n}/I_{n+1})}{n^{d-1}}$
is finite for ${\mathbb{Q}}$-divisorial filtrations of $m_{R}$-primary ideals, as is shown in Lemma 4.1. However, there are examples of graded families of ideals such that this limsup is infinite ([8, Theorem 4.3]). Even when this limsup is finite, it may be that the limsup is not a limit for graded filtrations of $m_{R}$-primary ideals ([8, Theorem 4.6]).
The function (3) is very nice in dimension 2. In Section 8 [15], it is shown that if $R$ is a two dimensional normal excellent local ring and =$\mathcal{I}=\{I_{n}\}$ is a ${\mathbb{Q}}$-divisorial filtration of $m_{R}$-primary ideals on $R$, then there exists a quadratic polynomial $P(n)$ with rational coefficients, a function $f(n)$ and a constant $c$ such that
@@
with $|f(n)|<c$ for all $n$. In [15, Theorem 9], it is shown that if $R$ has equicharacteristic zero then $f(n)$ is eventually periodic. Example 5 [15] shows that $f(n)$ may not be eventually periodic if $R$ has characteristic $p>0$.
@@
(5) $\lim_{n\rightarrow\infty}\frac{\lambda_{R}(I_{n}/I_{n+1})}{n^{d-1}}$
@@
As we have seen above, this limit does always exist and is even a rational number when the dimension of $R$ is $2$.
@@
It is shown in the formula before [10, Lemma 2.2] and [10, Theorem 4.1] that
@@
where $D_{n}=\lceil n\alpha\rceil\overline{S}+nF$, $\alpha=\frac{3}{9-\sqrt{3}}$. It is further shown that the divisors $-D_{n}$ are nef divisors on $U$. Thus by formula (9) and [10, Proposition 2.4],
@@ 1.4. Symbolic powers with infinitely many Rees valuations
Let $R$ be a $d$-dimensional excellent normal local ring. Let =$\mathcal{I}=\{I_{n}\}$ be a graded family of ideals of $R$ ($\mathcal{I}$ is a filtration of $R$ and $I_{m}I_{n}\subset I_{m+n}$ for all $m,n$). We have that $\sqrt{I_{n}}=\sqrt{I_{1}}$ for all $n$. We define the Rees valuations of $\mathcal{I}$ to be the divisorial valuations $v$ of $R$ such that the center of $v$ on $\mbox{Proj}(R[\mathcal{I}])$ has dimension $d-1$ and the center $\mbox{center}_{R}(v)$ of $v$ on $R$ contains $I_{1}$ (and so contains $I_{n}$ for all $n$). In the case that =$\mathcal{I}=\{I^{n}\}$ is the $I$-adic filtration for some ideal $I$ of $R$, this agrees with the classical definition of the Rees valuations of $I$ ([39, Chapter 10]).
Suppose that $\mathcal{I}$ is a divisorial filtration with $I_{n}=I(v_{1})_{n\lambda_{1}}\cap\cdots\cap I(v_{r})_{n\lambda_{r}}$ for all $n$. Then the Rees valuations of $\mathcal{I}$ are a subset of $\{v_{1},\ldots,v_{r}\}$. The Rees valuations can be a proper subset of $\{v_{1},\ldots,v_{r}\}$. Theorem 1.5 shows that if $R$ is essentially of finite type over a field (in addition to our assumption that $R$ is normal and excellent) then the analytic spread $\ell(\mathcal{I})=d$ if and only if there exists a Rees valuation $v$ of $\mathcal{I}$ such that $\mbox{center}_{R}(v)=m_{R}$.
We have the natural birational maps
$\mbox{Proj}(R[I_{n}t^{n}])\dashrightarrow\mbox{Proj}(R[\mathcal{I}])$
for all $n>0$. The nature of these birational maps (and hence the structure of $\mbox{Proj}(R[\mathcal{I}])$) is determined by the respective Rees valuations of $\mbox{Proj}(R[I_{n}t^{n}])$ and $\mbox{Proj}(R[\mathcal{I}])$.
This raises the question of what asymptotic properties the Rees valuations of $I_{n}$ have (as $n$ goes to infinity) and how this compares with the Rees valuations of $\mathcal{I}$.
If $I$ is an ideal in an excellent local ring, then we have seen in Theorem 1.3 that the filtration of powers of integral closures $\{\overline{I^{n}}\}$ of the powers of $I$ is a ${\mathbb{Z}}$-divisorial filtration. In particular, the set of all Rees valuations of all of the ideals $I^{m}$ over all $m$ is a finite set. It is just the set of Rees valuations of $I$ itself. In Section 5 we consider the question of if this set is finite for the ideals in a ${\mathbb{Z}}$-divisorial filtration; that is, is the set of all Rees valuations of the ideals $I_{n}$ finite when =$\mathcal{I}=\{I_{n}\}$ is a ${\mathbb{Z}}$-divisorial filtration? We show that this is not the case. In Section 5 we construct an example of a height one prime ideal $P$ in a two dimensional normal local ring $R$ such that for all $n\geq 1$, the symbolic power $P^{(n)}$ has a Rees valuation other then the $PR_{P}$-adic valuation $v_{P}$, but for $1\leq m<n$, $v_{P}$ is the only common Rees valuation of $P^{(m)}$ and $P^{(n)}$. In particular, the set of Rees valuations of $P^{(n)}$ over all $n\geq 1$ is infinite.
The only Rees valuation of the divisorial filtration $\{P^{(n)}\}$ is $v_{P}$ and the analytic spread of this filtration is ($\ell(\{P^{(n)}\})=0$. In contrast, the analytic spread of each individual ideal $P^{(n)}$ is maximal, with $\ell(P^{(n)})=2$. None of the Rees valuations of the $P^{(n)}$ which have center $m_{R}$ on $R$ have a center on $\mbox{Proj}(R[\mathcal{I}])=\mbox{Spec}(R)\setminus\{m_{R}\}$. Necessarily, $R[\mathcal{I}]$ is not Noetherian.
@@ 1.5. Acknowledgements
The author would like to thank Hailong Dao and Jonathan Montaño for discussions and questions about Hilbert polynomials of filtrations and to thank Linquin Ma, Thomas Polstra and Karl Schwede for a discussion about the cardinality (finite or infinite) of the set of Rees valuations of the symbolic powers of a prime ideal. He thanks Suprajo Das for suggesting the use of alterations to generalize the assumption that rings have equicharacteristic zero in [12, Theorem 1.3]. He thanks the reviewer for helpful comments to improve the exposition.